When copper metal is added to nitric acid, the following reaction takes place Cu (s) + 4 HNO₃ (aq) → Cu(NO₃)₂ (aq) + 2 H₂O (l) + 2 NO₂ (g) Calculate the volume of NO₂ gas collected over water at 25.0 °C when 8.13 g of copper is added to excess nitric acid if the total pressure is 726 mm Hg. The vapor pressure of water at 25.0 °C is 23.8 mm Hg.
Answer #1


Molar mass of Cu = 63.55 g/mol

mass of Cu = 8.13 g
mol of Cu = (mass)/(molar mass)
= 8.13/63.55
= 0.1279 mol

According to balanced equation
mol of NO2 formed = (2/1)* moles of Cu
= (2/1)*0.1279
= 0.2559 mol

Now for NO2:
P = 726 mmHg - 23.8 mmHg
= 702.2 mm Hg
= (702.2/760) atm
= 0.9239 atm
n = 0.2559 mol
T = 25.0 oC
= (25.0+273) K
= 298 K

use:
P * V = n*R*T
0.9239 atm * V = 0.2559 mol* 0.08206 atm.L/mol.K * 298 K
V = 6.7728 L
Answer: 6.77 L